Transverse position in Meld’s experiment is chanted to parallel position. If length of string remaining the same and tension is made half. If in perpendicular position 4 loops are formed number of loops formed in parallel position
(a) 1
(b) 2
(c) 3
(d) 4

Grade:Upto college level

2 Answers

Arun
3 years ago
Given :
p = 6  
L = 7.2 m
M = 10 g = 10 × 10⁻³ = 10⁻² kg
M′ = 14.4 × 10⁻² g = 14.4 × 10⁻⁵ kg
n = P / 2L√(T/ m)
For parallel position , frequency is given by  
N=2n
N= 2. P / 2L√(T/ m)
=P/L √(T/m)------1
As we know that m= M’/l
M=14.4x10-5/7.2
M=2x10-5 kg/m
Let us substitute this value in equation 1
N=6/7.2√ [ 10-2x9.8/2x10⁻⁶
=6/7.2 √[49x10⁻²]
=6x7x10/7.2
=70/1.2
=58.33Hz
∴ the frequency of tuning fork. is 58.33 Hz
Vikas TU
3 years ago
n = P / 2L√(T/ m)
For parallel position , frequency is given by  
N=2n
N= 2. P / 2L√(T/ m)
=P/L √(T/m)------1
As we know that m= M’/l
M=14.4x10-5/7.2
M=2x10-5 kg/m
Let us substitute this value in equation 1
N=6/7.2√ [ 10-2x9.8/2x10⁻⁶
=6/7.2 √[49x10⁻²]
=6x7x10/7.2
=70/1.2
=58.33Hz
∴ the frequency of tuning fork. is 58.33 Hz